DETERMINATION OF THE DIPOLE MOMENT OF H_{3}N-SO_{3}

\underline{MANJULA CANAGARATNA} , M.E. OTT, AND K.R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

The dipole moment of H_{3}N-SO_{3} has been determined using pulsed nozzle Fourier transform microwave spectroscopy. The value obtained, 6.204(12) D, is in reasonable agreement with published ab initio estimates ^{a}. These measurements follow our previously reported ^{b} structural work on this complex in which the intermolecular N-S bond was shown to be significantly longer in the gas phase than in the solid state. Presumably, the mechanism for the contraction of the N-S bond upon crystallization involves a sharply rising dipole moment function at shorter distances. Thus, the present result is especially interesting in connection with the observed gas-solid structure changes. The complex has also been postulated as a potential seed molecule for the nucleation of atmospheric aerosols, and we comment on possible implications of its phase-dependent structure and bonding on the kinetics and thermodynamics of the nucleation process.

aM. W. Wong, K. B. Wiberg, and M. J. Frisch, J. Am Chem. Soc. 114, 523 (1992).

bM. Canagaratna, J. A. Phillips, H. Goodfriend, and K. R. Leopold, J. Am. Chem. Soc., 118, 5290 (1996).